OpenStudy (anonymous):

Solve by completing the square: x2 + 8x = -3.

OpenStudy (anonymous):

first club on all the terms on LHS. so we get x2+8x+3=0. now add 16 and subtract 16. so equation will look like this x2+8x+16-3-16=0. now (x2+8x+16) can be shortened as (x+4)^2 so we get (x+4)^2-3-16=0 now you modify it and write the equation as (x+4)^2=19 now i hope you can solve this

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."